Suppose that it is given to you that
f′(x)=(x+6)(12−x)(x−15)

Then the first relative extremum (from the left) for f(x) occurs at x=

The function f(x) has a relative
?
at this point.

The second relative extremum (from the left) for f(x) occurs at x=

The function f(x) has a relative
?
at this point.

The third relative extremum (from the left) for f(x) occurs at x=

The function f(x) has a relative
?
at this point.

The first inflection point (from the left) for f(x) occurs at x=


The second inflection point (from the left) for f(x) occurs at x=
1
Expert's answer
2020-12-02T13:13:00-0500

Given Derivative is


"f'(x)=(x+6)(12-x)(x-15)"


For calculating critical points putting "f'(x)=0"


critical points are "-6,12,15" respectively.


The possible intervals are "(-\infty,-6),(-6,12),(12,15) ,(15,\infin)"


Sign of "f'(x) in (-\infty,-6)" is "(+)" "\implies f'(x)>0\implies f(x)" is increasing.



Sign of "f'(x) in (-6,12)" is "(-)" "\implies f'(x)<0\implies f(x)" is Decreasing.



Sign of "f'(x) in (12,15)" is "(+)" "\implies f'(x)>0\implies f(x)" is increasing.



Sign of "f'(x) in (15,\infty)" is "(-)" "\implies f'(x)<0\implies f(x)" is decreasing.


"\Rightarrow" The first relative extremum (from the left) for "f(x)" occurs at "x=-6"

The function "f(x)" has a relative maximum at this point.


"\Rightarrow" The second relative extremum (from the left) for "f(x)" occurs at "x=12"

The function "f(x)" has a relative minimum at this point.


"\Rightarrow" The third relative extremum (from the left) for "f(x)" occurs at "x=15"

The function "f(x)" has a relative maximum at this point.


There is no inflexion point for the given derivative.

Since the function changes its value for every interval.